The arithmetic mean (or average) of A, B and C is 10. The value of A is six less than the value of B, and the value of C is three more than the value of B. What is the value of C?
Answer by rfer(16322)

((b-6)+b+(b+3))/3=10
(3b-3)/3=10
3b-3=30
3b=33
b=11
a=5
c=14